OpenStudy (anonymous):

Which set of polar coordinates describes the same location as the rectangular coordinates (0,-2)

OpenStudy (paki):

the relation between polar coordinate n cartesian coordinate is \[r = \sqrt{{x^2 + y^2}}\]

OpenStudy (paki):

r = sqrt(x^2 + y^2) r = sqrt((0)^2 + (-2)^2).... solve for "r" here... and then use the relation... tan teeta = y/x...